# Self-Managing DRAM: A Low-Cost Framework for Enabling Autonomous and Efficient in-DRAM Operations

### v5

@@ 8 Related Work
To our knowledge, this is the first paper to propose simple changes to existing DRAM interfaces for enabling DRAM chips that autonomously perform various maintenance operations. Our new Self-Managing DRAM (SMD) design enables implementing maintenance operations completely within DRAM without requiring further changes in the DRAM interface, Memory Controller (MC) or other system components. We briefly discuss relevant prior works.
Changing the DRAM Interface. Several prior works [148, 37, 30, 22] propose using high-speed serial links and packed-based protocols in DRAM chips. SMD differs from prior works in two key aspects. First, none of these works describe how to implement maintenance mechanisms completely within DRAM. We propose five new SMD-based maintenance mechanisms (section 5). Second, prior works significantly overhaul the existing DRAM interface, which makes their proposals more difficult to adopt compared to SMD, which adds only a single ACT_NACK signal to the existing DRAM interface and requires slight modifications in the MC.
Mitigating DRAM Refresh Overhead. Many previous works [16, 86, 124, 107, 6, 8, 23, 26, 34, 46, 55, 63, 90, 64, 87, 99, 106, 117, 145, 57, 58, 59, 150, 121, 125, 39, 68, 110, 74] propose techniques to reduce DRAM refresh overhead. SMD not only enables maintenance mechanisms for reducing DRAM refresh overhead but also other efficient maintenance mechanisms for improving DRAM reliability and security, described in section 5.
RowHammer Protection. Many prior works [82, 139, 143, 155, 116, 159, 138, 5, 20, 70, 31, 43, 149, 12, 73, 70, 155, 123, 28, 67, 69, 7] propose techniques for RowHammer protection. Although we propose SMD-PRP, SMD-PRP+, and SMD-DRP, SMD can be used to implement other existing or new RowHammer protection mechanisms.
Memory Scrubbing. Although prior works report that the overhead of memory scrubbing is small as low scrubbing rate (e.g., scrubbing period of 24 hours [141, 126, 53], 45 minutes per 1 GB scrubbing rate [132], scrubbing only when idle [94]) is generally sufficient, the cost of scrubbing can dramatically increase for future DRAM chips due to increasing DRAM bit error rate and increasing DRAM chip density. SMD-MS enables efficient scrubbing by eliminating off-chip data transfers.
Leveraging Subarray-level Parallelism. Prior works overlap the latency of accessing multiple subarrays via modifications to the DRAM architecture [71, 161, 16]. SMD’s maintenance-access parallelization builds on the basic design proposed in SALP [71] and refresh-access parallelization introduced in [16, 161]. A DDRx interface based on these prior works [71, 161, 16] still needs to change for each new maintenance operation that the DRAM manufacturer needs to implement. SMD, in contrast, allows the implementation of new maintenance mechanisms without making any further changes to the DDRx interface.
@@ 9 Conclusion
Acknowledgments
We thank the SAFARI Research Group members for valuable feedback and the stimulating intellectual environment they provide. We acknowledge the generous gifts provided by our industrial partners, including Google, Huawei, Intel, Microsoft, and VMware, and support from the Microsoft Swiss Joint Research Center.

### v6

@@ 9 Related Work
This is the first paper to enable DRAM chips that autonomously and efficiently perform various maintenance operations with simple changes to existing DRAM interfaces. No prior work proposes setting the MC free from managing DRAM maintenance operations nor studies the system-level performance and energy impact of autonomous maintenance mechanisms. We briefly discuss relevant prior works.
Changing the DRAM Interface. Several prior works [129, 130, 131, 132] propose using high-speed serial links and packet-based protocols in DRAM chips. SMD differs from prior works in two key aspects. First, none of these works describe how to implement maintenance mechanisms completely within DRAM. We propose five new SMD-based maintenance mechanisms (section 5). Second, prior works significantly overhaul the existing DRAM interface, which makes their proposals more difficult to adopt compared to SMD, which adds only a single ACT_NACK signal to the existing DRAM interface and requires slight modifications in the MC.
Mitigating DRAM Refresh Overhead. Many previous works [16, 17, 15, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] propose techniques to reduce DRAM refresh overhead. SMD enables maintenance mechanisms for reducing DRAM refresh overhead and other efficient maintenance mechanisms for improving DRAM reliability and security (section 5).
RowHammer Protection. Many prior works [133, 49, 50, 54, 48, 53, 134, 45, 135, 10, 75, 136, 59, 58, 57, 137, 62, 63, 138, 139] propose techniques for RowHammer protection. One can use SMD to implement existing or new RowHammer protection mechanisms. Our goal is not to prove that SMD-DRP is the best RowHammer mitigation mechanism.
Memory Scrubbing. Although prior works report that the overhead of memory scrubbing is small as low scrubbing rate (e.g., scrubbing period of 24 hours [69, 70, 72], 45 minutes per 1 GB scrubbing rate [66], only when idle [67]) is generally sufficient, the cost of scrubbing can dramatically increase for future DRAM chips due to increasing DRAM bit error rate and increasing DRAM chip density. SMD-MS enables efficient scrubbing by eliminating off-chip data transfers.
Leveraging Subarray-level Parallelism. Prior works overlap the latency of accessing multiple subarrays by modifying the DRAM architecture [79, 80, 16]. SMD’s maintenance-access parallelization builds on the basic design proposed in SALP [79] and refresh-access parallelization introduced in [16, 80]. A DDRx interface based on these prior works [79, 80, 16] still needs to change for each new maintenance operation that the DRAM manufacturer needs to implement. SMD, in contrast, allows the implementation of new maintenance mechanisms without making any further changes to the DDRx interface.